# Optimal Multi-Distribution Learning

### v2

@@ 6.2 Algorithm and sample complexity
The proof of Lemma 15 is postponed to Appendix D
7 Extension: oracle-efficient multi-group learning
Given a hypothesis set $\mathcal{H}$ where each $h\in\mathcal{H}:\mathcal{X}\to\mathcal{Y}$, a set of groups $\mathcal{G}$ such that each $g\in\mathcal{G}$ is a subset of $\mathcal{X}$, and a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ and a distribution $\mathcal{D}$ with support $\mathcal{X}\times\mathcal{Y}$, define $L_{\mathcal{D}}(h|g):=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(h(x),y)|x\in g]$. Define $P_{g}=\mathrm{Pr}_{(x,y)\sim\mathcal{D}}[x\in g]$. Let $\gamma=\min_{g\in\mathcal{G}}P_{g}$. The goal of the learning algorithm is to find a (possibly randomized) hypothesis $h$ to minimize $L_{\mathcal{D}}(h|g)$ for all $g\in\mathcal{G}$ up to some threshold $\epsilon\in(0,1]$. That is,
$L_{\mathcal{D}}(h|g)\leq\min_{h^{\prime}\in\mathcal{H}}L_{\mathcal{D}}(h^{\prime}|g)+\epsilon,\forall g\in\mathcal{G}.$
We continue with the compatibility assumption as below, which ensures that such near-optimal hypothesis $h$ exists.
Assumption 16.
There exists $h^{*}\in\mathcal{H}$, such that
$L_{\mathcal{D}}(h^{*}|g)\leq\min_{h\in\mathcal{G}}L_{\mathcal{D}}(h|g)+\frac{\epsilon}{4},\quad\forall g\in\mathcal{G}.$
The learning algorithm
As presented in Algorithm 4, we first sample $N=O\left(\frac{\log(|\mathcal{G}|/\delta)+d\log(d/\epsilon)}{\gamma\epsilon^{2}}\right)$ datapoints from $\mathcal{D}$ and then estimate the optimal value $f_{g}$ for each group $g\in\mathcal{G}$. By taking $L_{\mathcal{D}}(h|g)-f_{g}$ as the loss function, we then invoke a Hedge algorithm over $\mathcal{G}$ to learn the problem
$\min_{h\in\Delta(\mathcal{H})}\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h|g)-f_{g}).$ (49)
By Assumption 16, it suffices to find some (possibly randomized) $h$ such that
$\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h|g)-f_{g}).\leq\min_{h^{\prime}\in\mathcal{H}}\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h^{\prime}|g)-f_{g})\leq\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h^{*}|g)-f_{g})=O(\epsilon).$
Formally, we have the following theorem.
Theorem 17.
Suppose Assumption 16 holds. By running Algorithm 4, with probability, the output $h^{\mathsf{final}}$ satisfies that
$L_{\mathcal{D}}(h^{\mathsf{final}}|g)=\frac{1}{T}\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)\leq\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)+\epsilon$ (50)
for any $g\in\mathcal{G}$. Meanwhile, the sample complexity of Algorithm 4 is bounded by
$O\left(\frac{d\log(d/\epsilon)+\log(|\mathcal{G}|/\delta)}{\gamma\epsilon^{2}}\right).$
Moreover, Algorithm 4 only access $\mathcal{H}$ with an $\mathrm{ERM}$ oracle.
Under Assumption 16, Theorem 17 recovers the result in Tosh and Hsu, (2022) without direct access to the hypothesis set $\mathcal{H}$. In high-level idea, the proof of Theorem 17 is based on the uniform convergence argument, which is a natural extension of our main results. We postpone the proof to Appendix E due to space limitation.
8 Necessity of randomization
Given that the best-known sample complexities prior to our work were derived for algorithms that either output randomized hypotheses or invoke majority votes, Awasthi et al., (2023) raised the question about how the sample complexity is impacted if only deterministic (or “proper”) hypotheses are permitted as the output of the learning algorithms. As it turns out, the restriction to deterministic hypotheses substantially worsens the sample efficiency, as revealed by the following theorem.
Theorem 18.
  • a hypothesis class $\mathcal{H}$ containing at most $kN_{0}+1$ hypothesis,
such that it takes at least $\frac{dk}{240000\varepsilon^{2}}$ samples to find $h\in\mathcal{H}$ obeying
with probability exceeding $3/4$.
Let us breifly desribe the high-level strategy for our construction of the hard instance: for each $i\in[k]$, we build a hypothesis set $\mathcal{H}_{i}$ that performs poorly solely on $\mathcal{D}_{i}$. To discriminate the optimal hypothesis — denoted by $h^{\star}$ — from some $\mathcal{H}_{i}$, the learner has to call $\Omega(\log(|\mathcal{H}_{i}|)/\varepsilon^{2})=\widetilde{\Omega}(d/\varepsilon^{2})$ times to $\mathrm{Query}(\mathcal{D}_{i})$ in expectation. The result follows by taking sum over $i\in[k]$.
Proof of Theorem 18.
Note that $N_{0}=\frac{2^{d}-1}{k}$. Set $N=kN_{0}+1=2^{d}$. Set $\mathcal{X}=\{-1,0,1\}^{kN}$. We set $\mathcal{Y}=\{1\}$ to be a set with only one element. Without loss of generality we write $l(h,(x,y))=l(h,x)$.
We now introduce our construction. There are $N$ hypotheses in $\mathcal{H}$, where each hypothesis corresponds to $k$ dimensions of the ground set $\mathcal{X}$. Without loss of generality, for $h\in\mathcal{H}$, we let $\mathcal{I}_{h}=\{j_{h,i}\}_{i=1}^{k}$ be the $k$ dimensions related to $h$. Note that $\mathcal{I}_{h}\cap\mathcal{I}_{h^{\prime}}=\emptyset$ for $h\neq h^{\prime}$. We construct $\mathcal{H}$ as $\mathcal{H}=(\cup_{i=1}^{k}\mathcal{H}_{i})\cup\{h^{\star}\}$. Then we define $l(h,x)$ as $l(h,x)=\min_{i\in\mathcal{I}_{h},x_{i}\neq 0}x_{i}$. Now we design the $k$ distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$. Fix $i\in[k]$, we let $\mathrm{Pr}_{\mathcal{D}_{i}}[x]=\Pi_{\ell=1}^{kN}\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]$, where $\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]=\mathbb{I}[x_{\ell}=0]$ for $\ell\notin\{j_{h,i}|h\in\mathcal{H}\}$, $\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]=\frac{1}{2}\mathbb{I}[x_{\ell}=1]+\frac{1}{2}\mathbb{I}[x_{\ell}=-1]$ for $\ell\in\{j_{h,i}|h\notin\mathcal{H}_{i}\}$, and $\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]=\left(\frac{1}{2}+4\varepsilon\right)\mathbb{I}[x_{\ell}=1]+\left(\frac{1}{2}-4\varepsilon\right)\mathbb{I}[x_{\ell}=-1]$ for $\ell\in\{j_{h,i}|h\in\mathcal{H}_{i}\}$. In words, for each $i\in[k]$, we draw $x\sim\mathcal{D}_{i}$ by independently draw each dimension of $x$, and there are $|\mathcal{H}|=N$ non-zero dimensions. Furthermore, for $x\sim\mathcal{D}_{i}$ and $h\in\mathcal{H}$, we have $l(h,x)=x_{j_{h,i}}$. By this construction, we have the following properties:
  1. (i).
    $\ell(h,x)\in[-1,1]$ for any $h\in\mathcal{H}$ and $x\in\mathcal{X}$;
  2. (ii).
    $\mathbb{E}_{x\sim\mathcal{D}_{i}}[\ell(h,x)]=\mathbb{E}_{x\sim\mathcal{D}_{i,j_{h,i}}}[x_{j_{h,i}}]=2\varepsilon\cdot\mathbb{I}[h\in\mathcal{H}_{i}]$ for any $i\in[k]$ and $h\in\mathcal{H}$;
  3. (iii).
    the only $\varepsilon$-optimal hypothesis is $h^{\star}$ because for any $h\neq h^{\star}$, there exists some $i$ such that $h\in\mathcal{H}_{i}$;
  4. (iv).
    $|\mathcal{H}|=N=kN_{0}+1=2^{d}$, which implies $\mathrm{VC}(\mathcal{H})\leq\log_{2}(N)\leq d$.
For each call to $\mathrm{Query}(\mathcal{D}_{i})$, we can get independent observations $\{x_{j_{h,i}}\}_{h\in\mathcal{H}}$ where $x_{j_{h,i}}\sim D_{i,j_{h,i}}$ for each $h\in\mathcal{H}$.
Now we let the number of calls to $\mathrm{Query}(\mathcal{D}_{i})$ be $M_{i}$ for $i\in[k]$. Our target is to show to distinguish $h^{\star}$ from $\mathcal{H}_{i}$, $M_{i}$ has to be at least $\Omega(d/\varepsilon^{2})$.
Now suppose there is an algorithm $\mathcal{G}$ with numbers of samples $\{M_{i}\}_{i=1}^{k}$ such that the output is $h^{\star}$ with probability $\frac{3}{4}$. Let $\mathrm{Pr}_{\mathcal{G}}[\cdot]$ and $\mathbb{E}_{\mathcal{G}}[\cdot]$ denote respectively the probability and expectation under running the algorithm $\mathcal{G}$. Let $h_{\mathrm{out}}$ be the output hypothesis. It then follows that
$\mathrm{Pr}_{\mathcal{G}}[h_{\mathrm{out}}=h^{\star}]\geq\frac{3}{4}.$
Let $\Pi_{\mathcal{H}}$ be the set of permutations over $\mathcal{H}$. Let $\mathrm{U}(\Pi_{\mathcal{H}})$ be the uniform distribution over $\Pi(\mathcal{H})$. With a slight abuse of notations, for $x\in\{-1,0,1\}^{kN}$ and $\sigma\in\Pi_{\mathcal{H}}$, we define $\sigma(x)$ to be the vector $y$ such that $y_{j_{h,i}}=x_{j_{\sigma(h),i}}$ for all $h\in\mathcal{H}$ and $i\in[k]$.
Let $\mathcal{G}^{\prime}$ be the algorithm with $\mathcal{H}$ replaced by $\sigma(\mathcal{H})$ in the input where $\sigma\sim\mathrm{U}(\Pi_{\mathcal{H}})$. Because $\mathcal{G}$ returns the optimal hypothesis with probability at least $3/4$ for all problem instances, we have that
$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star}]\geq\frac{3}{4}.$
We note that $\mathcal{G}^{\prime}$ is a symmetric algorithm with respect to the hypothesis set. Formally, we have the lemma below to bound the probability of outputting a non-optimal hypothesis.
Lemma 19.
Fix $m\geq 0$ and $\widetilde{i}\in[k]$. Suppose $\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{i}\leq m]\geq\frac{1}{2}$. It then holds that for any $h\in\mathcal{H}_{\widetilde{i}}$
$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m]\geq\frac{1}{2}\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m]\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}).$
Moreover, it holds that $m\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}$.
By Lemma 19, we have that
@@ 8 Necessity of randomization
$\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right]\leq\frac{1}{2}.$ (52)
Noting that $\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star}]\geq\frac{3}{4}$, we learn that
$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}]\geq\frac{1}{4},$ (53)
which implies that $\mathbb{E}_{\mathcal{G}^{\prime}}[M_{\widetilde{i}}]\geq\frac{\log(N_{0}/4)}{120000\varepsilon^{2}}\geq\frac{d-\log_{2}(8k)}{120000\varepsilon^{2}}\geq\frac{d}{240000\varepsilon^{2}}$. Taking sum over $i\in[k]$, we have that
$\mathbb{E}_{\mathcal{G}^{\prime}}\left[\sum_{i=1}^{k}M_{\widetilde{i}}\right]\geq\frac{dk}{240000\varepsilon^{2}}.$ (54)
The proof is completed.
@@ 9 Discussion
In this paper, we have settled the problem of achieving optimal sample complexity in multi-distribution learning, assuming the availability of adaptive (or on-demand) sampling. We have put forward a novel oracle-efficient algorithm that provably attains a sample complexity of $\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$ for VC classes, which matches the best-known lower bound up to some logarithmic factor. From the technical perspective, the key novelty of our analysis lies in carefully bounding the trajectory of the Hedge algorithm on a convex (concave) optimization problem. We have further unveiled the necessesity of randomization, revealing that a considerably larger sample size is necessary if the learning algorithm is constrained to return deterministic hypotheses. Notably, our work manages to solve three open problems presented in COLT 2023 (namely, Awasthi et al., (2023, Problems 1, 3 and 4)).
Our work not only addresses existing challenges but also opens up several directions for future exploration. To begin with, while our sample complexity results are optimal up to logarithmic factors, further studies are needed in order to sharpen the logarithmic depdency. Additionally, the current paper assumes a flexible sampling protocol that allows the learner to take samples arbitrarily from any of the $k$ distributions; how will the sample complexity be impacted under additional constraints imposed on the sampling process? Furthermore, can we extend our current analysis (which bounds the dynamics of the Hedge algorithm) to control the trajectory of more general first-order/second-order algorithms, in the context of robust online learning? Another venue for exploration is the extension of our multi-distribution learning framework to tackle other related tasks like multi-calibration (Hébert-Johnson et al.,, 2018; Haghtalab et al.,, 2023). We believe that our algorithmic and analysis framework can shed light on making progress in all of these directions.
@@ Appendix A Auxiliary lemmas
We begin by introducing three handy concentrations inequalities. The first result is the well-renowned Freedman inequality (Freedman,, 1975), which assists in deriving variance-aware concentration inequalities for martingales.
Lemma 20 (Freedman’s inequality (Freedman,, 1975)).
@@
Lemma 21 (Lemma 10 in Zhang et al., (2022)).
@@
Appendix B Missing proofs for optimality
@@ B.1 Proof of Lemma 4
Consider any fixed $n=\{n_{i}\}_{i=1}^{k}$ obeying $0\leq n_{i}\leq T_{1}$ for all $i\in[k]$, and any fixed $w\in\Delta(k)$. For any quantity $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$, if we take
$E(\lambda)\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right\}\right)\right]$ (57)
with the expectation taken over the randomness of $\widetilde{\mathcal{S}}(n)$, then we can perform a standard “symmetrization” trick to bound $E(\lambda)$ as follows:
$E(\lambda)$ $\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right\}\right)\right]$
@@
Note that for any dataset $\mathcal{C}$ with cardinality $|\mathcal{C}|$, the Sauer–Shelah lemma (Wainwright,, 2019, Proposition 4.18) together with our assumption that $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$ tells us that the cardinality of the following set obeys
$\big{|}\mathcal{H}(\mathcal{C})\big{|}\leq(|\mathcal{C}|+1)^{d}\leq(2kT_{1}+1)^{d},$ (60)
@@
$\mathop{\mathbb{E}}\limits_{\sigma_{i,j}}\big{[}\exp(\sigma_{i,j}x)\big{]}=\frac{1}{2}\big{(}\exp(x)+\exp(-x)\big{)}\leq\exp\big{(}0.5x^{2}\big{)}.$
Taking (58) and (62) together and applying the Markov inequality reveal that, for any $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$,
$E(\lambda)\leq(2\|n\|_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right).$ (63)
Repeating the same arguments also yields
$\overline{E}(\lambda)$ $\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}w_{i}L(h,e_{i})-\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}\right\}\right)\right]$
@@
for any $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$. Taking the above two inequalities reveals that, for any $0<\varepsilon^{\prime}\leq 1$,
$\mathbb{P}\left(\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right|\geq\varepsilon^{\prime}\right)$
@@
Step 2: uniform concentration bounds over all $n$ and $w$.
Next, we move on to extend the above result to uniform concentration bounds. Towards this, let us first introduce a couple of notation.
@@
    =$\mathcal{B}=\Big{\{}n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\mid T_{1}w_{i}\leq 2n_{i},0\leq n_{i}\leq T_{1},\forall i\in[k],w\in\Delta_{\varepsilon_{1}/(8k)}(k)\Big{\}}.$
Applying the union bound yields
$\mathbb{P}\left(\exists(n,w)\in\mathcal{B},\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right|\geq\varepsilon^{\prime}\right)$
@@
$\leq\sum_{(n,w)\in\mathcal{B}}\min_{0\leq\lambda\leq\frac{T_{1}}{2}}2\cdot(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\frac{2}{T_{1}}-\lambda\varepsilon^{\prime}\right)$
@@
$\leq|\mathcal{B}|\cdot 2\cdot(2kT_{1}+1)^{d}\exp\left(\frac{-T_{1}(\varepsilon^{\prime})^{2}}{16}\right)$
$\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(\frac{-T_{1}(\varepsilon^{\prime})^{2}}{16}\right),$

### v3

@@ 10                   $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
The proof of Lemma 6 is postponed to Appendix E. Combine this with Lemma 2 and Lemma 3 to show that: the total number of samples collected in Algorithm 3 is upper bounded by
@@
$T_{1}\,\mathrm{poly}\log\Big{(}k,\frac{1}{\varepsilon},\frac{1}{\delta}\Big{)}\leq\left(\frac{k}{\varepsilon^{2}}+\min\bigg{\{}C_{n}\mid C_{n}\leq\frac{1}{4800}\varepsilon\bigg{\}}\right)\,\mathrm{poly}\log\Big{(}k,\frac{1}{\varepsilon},\frac{1}{\delta}\Big{)}$
as claimed. ∎
@@ 8 Discussion
In this paper, we have settled the problem of achieving optimal sample complexity in multi-distribution learning, assuming availability of adaptive (or on-demand) sampling. We have put forward a novel oracle-efficient algorithm that provably attains a sample complexity of $\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$ for VC classes, which matches the best-known lower bound up to some logarithmic factor. From the technical perspective, the key novelty of our analysis lies in carefully bounding the trajectory of the Hedge algorithm on a convex-concave optimization problem. We have further unveiled the necessesity of randomization, revealing that a considerably larger sample size is necessary if the learning algorithm is constrained to return deterministic hypotheses. Notably, our work manages to solve three open problems presented in COLT 2023 (namely, awasthi2023sample).
Our work not only addresses existing challenges but also opens up several directions for future exploration. To begin with, while our sample complexity results are optimal up to logarithmic factors, further studies are needed in order to sharpen the logarithmic depdency. Additionally, the current paper assumes a flexible sampling protocol that allows the learner to take samples arbitrarily from any of the $k$ distributions; how will the sample complexity be impacted under additional constraints imposed on the sampling process? Furthermore, can we extend our current analysis (which bounds the dynamics of the Hedge algorithm) to control the trajectory of more general first-order/second-order algorithms, in the context of robust online learning? Another venue for exploration is the extension of our multi-distribution learning framework to tackle other related tasks like multi-calibration (hebert2018multicalibration; haghtalab2023unifying). We believe that our algorithmic and analysis framework can shed light on making progress in all of these directions.
@@ Appendix A Auxiliary lemmas
We begin by introducing three handy concentrations inequalities. The first result is the well-renowned Freedman inequality (freedman1975tail), which assists in deriving variance-aware concentration inequalities for martingales.
Lemma 7 (Freedman’s inequality (freedman1975tail)).
@@
Lemma 8 (Lemma 10 in zhang2022horizon).
@@
Finally, let us present a basic property related to Rademacher random variables, which will play a useful role in understanding the Rademacher complexity.
Lemma 12.
Let $\mathcal{L}$ be a set of vectors in $\mathbb{R}^{n}$. Let $w^{1},w^{2}\in\mathbb{R}^{n}$ be two vectors obeying $|w^{1}_{i}|\leq|w_{i}^{2}|$ for all $i\in[n]$. Then it holds that
$\mathop{\mathbb{E}}_{\{\sigma_{i}\}_{i=1}^{n}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w^{1}_{i}f_{i}\right]\leq\mathop{\mathbb{E}}_{\{\sigma_{i}\}_{i=1}^{n}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w^{2}_{i}f_{i}\right],$ (44)
where $\{\sigma_{i}\}$ is a collection of independent Rademacher random variables obeying $\mathbb{P}(\sigma_{i}=1)=\mathbb{P}(\sigma_{i}=-1)=1/2$.
Proof.
Clearly, it suffices to prove (44) for the special case where $w_{i}^{1}=w^{2}_{i}$ for $11\leq i\leq n-$ 1, and $|w^{1}_{n}|\leq|w^{2}_{n}|$. Fixing $\sigma_{i}$ for $11\leq i\leq n-$ 1, we can deduce that
$\mathop{\mathbb{E}}_{\sigma_{n}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w_{i}f_{i}\right]$ $=\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}f_{i}+w_{n}f_{n}\right)+\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}f_{i}-w_{n}f_{n}\right)$
$=\frac{1}{2}\max_{f,\widetilde{f}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}(f_{i}+\widetilde{f}_{i})+w_{n}(f_{n}-\widetilde{f}_{n})\right)$
$=\frac{1}{2}\max_{f,\widetilde{f}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}\widetilde{w}_{i}(f_{i}+\widetilde{f}_{i})+w_{n}(f_{n}-\widetilde{f}_{n})\right)$
$\leq\frac{1}{2}\max_{f,\widetilde{f}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}\widetilde{w}_{i}(f_{i}+\widetilde{f}_{i})+\big{|}\widetilde{w}_{n}(f_{n}-\widetilde{f}_{n})\big{|}\right)$
$=\frac{1}{2}\max_{f,\widetilde{f}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}\widetilde{w}_{i}(f_{i}+\widetilde{f}_{i})+\widetilde{w}_{n}(f_{n}-\widetilde{f}_{n})\right)$
$=\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}\widetilde{w}_{i}f_{i}+\widetilde{w}_{n}f_{n}\right)+\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}\widetilde{w}_{i}f_{i}-\widetilde{w}_{n}f_{n}\right)$
$=\mathop{\mathbb{E}}_{\sigma_{n}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}\widetilde{w}_{i}f_{i}\right].$
The proof is thus completed by taking expectation over $\{\sigma_{i}\}_{i=1}^{n-1}$. ∎
Appendix B Proofs of auxiliary lemmas for VC classes
@@ B.1 Proof of Lemma 1
Consider first any fixed $n=\{n_{i}\}_{i=1}^{k}$ obeying $0\leq n_{i}\leq T_{1}$ for all $i\in[k]$, and any fixed $w\in\Delta(k)$. For any quantity $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$, if we take
$E(\lambda,n,w)\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,w)\right\}\right)\right]$ (45)
with the expectation taken over the randomness of $\widetilde{\mathcal{S}}(n)$, then we can apply a standard “symmetrization” trick to bound $E(\lambda,n,w)$ as follows:
$E(\lambda,n,w)$ $\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,w)\right\}\right)\right]$
@@
Note that for any dataset $\mathcal{C}$ with cardinality $|\mathcal{C}|$, the Sauer–Shelah lemma (wainwright2019high, Proposition 4.18) together with our assumption that $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$ tells us that the cardinality of the following set obeys
$\big{|}\mathcal{H}(\mathcal{C})\big{|}\leq(|\mathcal{C}|+1)^{d}\leq\big{(}|\widetilde{\mathcal{S}}|+|\widetilde{\mathcal{S}}^{+}|+1\big{)}^{d}\leq(2kT_{1}+1)^{d},$ (48)
@@
$\mathop{\mathbb{E}}\limits_{\sigma_{i,j}}\big{[}\exp(\sigma_{i,j}x)\big{]}=\frac{1}{2}\big{(}\exp(x)+\exp(-x)\big{)}\leq\exp\big{(}x^{2}\big{)}.$
Taking (46), (47) and (50) together reveals that
$E(\lambda)\leq(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right).$ (51)
Repeating the same arguments also yields an upper bound on the following quantity:
$\overline{E}(\lambda)$ $\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{L(h,w)-\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}\right\}\right)\right]$
@@
for any $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$. Taking the above two inequalities and applying the Markov inequality reveal that, for any $0<\varepsilon^{\prime}\leq 1$,
$\mathbb{P}\left(\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,w)\right|\geq\varepsilon^{\prime}\right)$
@@
Step 2: uniform concentration bounds over epsilon-nets w.r.t. $n$ and $w$.
Next, we move on to extend the above result to uniform concentration bounds over all possible $n$ and $w$. Towards this, let us first introduce a couple of notation.
@@
    =$\mathcal{B}=\bigg{\{}n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\,\Big{|}\,\frac{n_{i}}{w_{i}}\geq\frac{T_{1}}{2},0\leq n_{i}\leq T_{1},\forall i\in[k],w\in\Delta_{\varepsilon_{1}/(8k)}(k)\bigg{\}},$
    which clearly satisfies
    $|\mathcal{B}|\leq T_{1}^{k}\cdot\left(\frac{8k}{\varepsilon_{1}}\right)^{k}.$
Applying the union bound yields that, for any $0<\varepsilon^{\prime}\leq 1$,
$\mathbb{P}\left(\exists(n,w)\in\mathcal{B},\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,w)\right|\geq\varepsilon^{\prime}\right)$
@@
$\leq\sum_{(n,w)\in\mathcal{B}}\min_{0\leq\lambda\leq\frac{T_{1}}{2}}2\cdot(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\cdot\frac{2}{T_{1}}-\lambda\varepsilon^{\prime}\right)$
@@
$\leq|\mathcal{B}|\cdot 2\cdot(2kT_{1}+1)^{d}\exp\left(-\frac{T_{1}(\varepsilon^{\prime})^{2}}{16}\right)$
$\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(-\frac{T_{1}(\varepsilon^{\prime})^{2}}{16}\right),$